OpenStudy (anonymous):

6^2x8^2 can be written 6^2x8^2^3 why can thee exponent 1 be removed?

OpenStudy (jdoe0001):

exponent one ? I don't see that one

OpenStudy (anonymous):

ugh my bad well yeah thats what i said but i guess its a typo its 2^1 i think

OpenStudy (anonymous):

not 2^3

OpenStudy (jdoe0001):

did you mean \(\large 6^2\times 8^2 \implies 6^2\times 8^{2^{\color{blue}{ 1?}}}\) by any chance?

OpenStudy (anonymous):

yeah lol

OpenStudy (jdoe0001):

well.. any number raised to 1, is just itself, thus an exponent simply states how many times a number must add by itself so "2" raised to "1" means " add 2 by itself one time"

OpenStudy (anonymous):

so much work...

OpenStudy (anonymous):

oh so 4?

OpenStudy (jdoe0001):

4?

OpenStudy (anonymous):

you said add 2 by itself. can i just have the answer im so brain dead i just want to finish the class

OpenStudy (jdoe0001):

any number raised at 1, is just itself

OpenStudy (anonymous):

okay so how do i solve it?

OpenStudy (anonymous):

oh nvm would it be the same? idk like 2^6?

OpenStudy (jdoe0001):

hmmm what are you mean to solve again? looks just like a simple multiplcation

OpenStudy (anonymous):

its expanded form exponents

OpenStudy (anonymous):

and yeah it does but idk

OpenStudy (anonymous):

its a word question

OpenStudy (jdoe0001):

ohhh hahemm \(\large \bf 8\to 2\cdot 2\cdot 2\to 2^3\qquad thus\implies 8^2\implies (2^3)^2\) if that's what you mean

OpenStudy (anonymous):

yeeeeaahhh that looks like it could be right

OpenStudy (jdoe0001):

well... you asked "wordedly" if the 1 could be removed.... so hmm thus what I said

OpenStudy (jdoe0001):

well... the exponent states how many times the number must multiplly by itself, no add and the multiplication is an addition of it

OpenStudy (anonymous):

my teacher will understand it